OpenStudy (anonymous):

using-only numbers, and not letters, write your own equation that illustrate the distributive property your equation should show that you are `distributing`` multiplication over addition instead of adding what inside the parentheses first (10a)58=(58x10)a

OpenStudy (anonymous):

what you have there is a combination of associative property and commutative property of multiplication distribution would be like 3(1+2)= 3*1+3*2=3+6

OpenStudy (anonymous):

you need a plus sign inside the parentheses like for example \[5\times (2+8)\]